### v1

# Harnessing the Universal Geometry of Embeddings

Abstract
We introduce the first method for translating text embeddings from one vector space to another without any paired data, encoders, or predefined sets of matches. Our unsupervised approach translates any embedding to and from a universal latent representation (i.e., a universal semantic structure conjectured by the Platonic Representation Hypothesis). Our translations achieve high cosine similarity across model pairs with different architectures, parameter counts, and training datasets.
The ability to translate unknown embeddings into a different space while preserving their geometry has serious implications for the security of vector databases. An adversary with access only to embedding vectors can extract sensitive information about the underlying documents, sufficient for classification and attribute inference.
Figure 1: Left: input embeddings from different model families (T5-based GTR [41] and BERT-based GTE [29]) are fundamentally incomparable. Right: given unpaired embedding samples from different models on different texts, our model learns a latent representation where they are closely aligned.
Figure 2: Given only a vector database from an unknown model, vec2vec translates the database into the space of a known model using latent structure alone. Converted embeddings reveal sensitive information about the original documents, such as the topic of an email (pictured, real example).
1 Introduction
Text embeddings are the backbone of modern NLP, powering tasks like retrieval, RAG, classification, and clustering. There are many embedding models trained on different datasets, data shufflings, and initializations. An embedding of a text encodes its semantics: a good model maps texts with similar semantics to vectors close to each other in the embedding space. Since semantics is a property of text, different embeddings of the same text should encode the same semantics. In practice, however, different models encode texts into completely different and incompatible vector spaces.
The Platonic Representation Hypothesis [17] conjectures that all image models of sufficient size have the same latent representation. We propose a stronger, constructive version of this hypothesis for text models: the universal latent structure of text representations can be learned and, furthermore, harnessed to translate representations from one space to another without any paired data or encoders.
In this work, we show that the Strong Platonic Representation Hypothesis holds in practice. Given unpaired examples of embeddings from two models with different architectures and training data, our method learns a latent representation in which the embeddings are almost identical (Figure˜1).
We draw inspiration from research on aligning word embeddings across languages [54, 9, 14, 8] and unsupervised image translation [32, 60]. Our vec2vec method uses adversarial losses and cycle consistency to learn to encode embeddings into a shared latent space and decode with minimal loss. This makes unsupervised translation possible. We use a basic adversarial approach with vector space preservation [40] to learn a mapping from an unknown embedding distribution to a known one.
vec2vec is the first method to successfully translate embeddings from the space of one model to another without any paired data. vec2vec translations achieve cosine similarity as high as $0.92$ to the ground-truth vectors in their target embedding spaces and perfect matching on over $8000$ shuffled embeddings (without access to the set of possible matches in advance).
To show that our translations preserve not only the relative geometry of embeddings but also the semantics of underlying inputs, we extract information from them using zero-shot attribute inference and inversion, without any knowledge of the model that produced the original embeddings!
2 Problem formulation: unsupervised embedding translation
Consider a collection of embedding vectors $\{u_{1},\ldots u_{n}\}$, for example, a dump of a compromised vector database, where each $u_{i}=M_{1}(d_{i})$ is generated by an unknown encoder $M_{1}:\mathbb{V}^{s}\rightarrow\mathbb{R}^{d_{M_{1}}}$ from an unknown document $d_{i}$. We cannot make queries to $M_{1}$ and do not know its training data, nor architectural details. Our goal is to extract any information about the documents $d_{i}$.
Figure 3: Unsupervised embedding translation. With access to only $u_{i}=M_{1}(d_{i})$, vec2vec seeks to generate a translation $F(u_{i})$ that is close in $M_{2}$’s embedding space to the ideal embedding $v_{i}=M_{2}(d_{i})$ without access to $d_{i}$, $v_{i}$, $M_{1}$, or $M_{2}$.
We do assume access to a different encoder $M_{2}$ that we can query at will to generate new embeddings in some other space. We also assume high-level distributional knowledge about the hidden documents: their modality (text) and language (e.g., English). To extract information, we may translate $\{u_{1},\ldots u_{n}\}$ into the output space of $M_{2}$ and apply techniques like inversion that take advantage of the encoder.
Correspondence methods cannot be used. There is significant prior research on the problem of matching or correspondence between sets of embedding vectors [1, 43, 7, 48]. None of these approaches are applicable in our setting because they assume the existence of two (or more) sets of embedding vectors produced by different encoders from the same inputs. In other words, for each unknown vector, there must already exist a set of candidate vectors in a different embedding, one of which is the correct match. In practice, it is unrealistic to expect that such a database be available.  [^1] Our code is available on GitHub.
In our setting, where we do not assume access to the encoder $M_{1}$, nor any additional representations of the unknown documents $\{d_{1},\ldots,d_{n}\}$ other than their embeddings $u_{i}=M_{1}(d_{i})$, the problem is strictly harder and requires unsupervised translation from $M_{1}$ to $M_{2}$. Any success at unsupervised translation will come solely from the common geometric structure of $M_{1}$’s and $M_{2}$’s output spaces.
The Strong Platonic Representation Hypothesis. Our hope that unsupervised embedding translation is possible at all rests on the stronger version of the Platonic Representation Hypothesis [17]. Our conjecture is as follows: neural networks trained with the same objective and modality, but with different data and model architectures, converge to a universal latent space such that a translation between their respective representations can learned without any pairwise correspondence.
Translation enables information extraction. Solving unsupervised translation will allow us to use information extraction tools designed to operate on vectors produced by known encoders. For example, we could apply inversion models [37, 58] to recover unknown documents $\{d_{i}\}$.
3 Our method: vec2vec
Unsupervised translation has been successful in computer vision, using a combination of cycle consistency and adversarial regularization [32, 60]. Our design of vec2vec is inspired in part by these methods. We aim to learn embedding-space translations that are cycle-consistent (mapping to and from an embedding space should end in the same place) and indistinguishable (embeddings for the same text from either space should have identical latents).
3.1 Architecture
We propose a modular architecture, where embeddings are encoded and decoded using space-specific adapter modules and passed through a shared backbone network. Figure˜2 shows these components. Input adapters $A_{1}:\mathbb{R}^{d}\to\mathbb{R}^{Z}$ and $A_{2}:\mathbb{R}^{d}\to\mathbb{R}^{Z}$ transform embeddings from each encoder-specific space into a universal latent representation of dimension $Z$. The shared backbone $T:\mathbb{R}^{Z}\to\mathbb{R}^{Z}$ extracts a common latent embedding from adapted inputs. Output adapters $B_{1}:\mathbb{R}^{Z}\to\mathbb{R}^{d}$ and $B_{2}:\mathbb{R}^{Z}\to\mathbb{R}^{d}$ translate these common latent embeddings back into the encoder-specific spaces. Thus, translation functions $F_{1},F_{2}$ and additional reconstruction mappings $R_{1},R_{2}$ are defined as:
$F_{1}=B_{2}\circ T\circ A_{1},\quad F_{2}=B_{1}\circ T\circ A_{2}\quad R_{1}=B_{1}\circ T\circ A_{1}\quad R_{2}=B_{2}\circ T\circ A_{2}$
Parameters of all components are collectively denoted $\theta=\{A_{1},A_{2},T,B_{1},B_{2}\}$.
Unlike images, embeddings do not have any spatial bias. Instead of CNNs, we use multilayer perceptrons (MLP) with residual connections, layer normalization, and SiLU nonlinearities. Discriminators mirror this structure but omit residual connections to simplify adversarial learning.
3.2 Optimization
In addition to the ‘generator’ networks $F$ and $R$, we introduce discriminators operating on both the latent representations of $F$ ($D_{1}^{\ell},D_{2}^{\ell}$) and the output embeddings ($D_{1},D_{2}$).
Our goal is to train the parameters of $\theta$ by solving:
$\theta^{*}=\arg\min_{\theta}\max_{D_{1},D_{2},D_{1}^{\ell},D_{2}^{\ell}}\mathcal{L}_{\text{adv}}(F_{1},F_{2},D_{1},D_{2},D_{1}^{\ell},D_{2}^{\ell})+\lambda_{\text{gen}}\mathcal{L}_{\text{gen}}(\theta),$ (1)
where $\mathcal{L}_{\text{adv}}$ and $\mathcal{L}_{\text{gen}}$ represent adversarial and generator-specific constraints respectively and hyperparameter $\lambda_{\rm gen}$ controls their tradeoff.
Adversarial loss. The adversarial loss encourages generated embeddings to match the empirical distributions of original embeddings both at the embedding and latent levels. Specifically, applying the standard GAN loss formulation [12] to both levels yields:
$\mathcal{L}_{\text{adv}}(F_{1},F_{2},D_{1},D_{2},D_{1}^{\ell},D_{2}^{\ell})$ $=\mathcal{L}_{\text{GAN}}(D_{1},F_{1})+\mathcal{L}_{\text{GAN}}(D_{2},F_{2})$
$\quad+\mathcal{L}_{\text{GAN}}(D_{1}^{\ell},T\circ A_{1})+\mathcal{L}_{\text{GAN}}(D_{2}^{\ell},T\circ A_{2}).$
Generator. Because adversarial losses alone do not guarantee that translated embeddings preserve semantics [60], we introduce three additional constraints to help the generator learn a useful mapping:
Reconstruction enforces that an embedding, when mapped into the latent space and back into its original embedding space, closely matches its initial representation:
$\mathcal{L}_{\text{rec}}(R_{1},R_{2})=\mathbb{E}_{x\sim p}\|R_{1}(x)-x\|_{2}^{2}+\mathbb{E}_{y\sim q}\|R_{2}(y)-y\|_{2}^{2}.$
where $p$ and $q$ are distributions of embeddings sampled from $M_{1}$ and $M_{2}$, respectively.
Cycle-consistency acts as an unsupervised proxy for supervised pair alignment, ensuring that $F$ and $G$ can translate an embedding to the other embedding space and back again with minimal corruption:
$\mathcal{L}_{\text{CC}}(F_{1},F_{2})=\mathbb{E}_{x\sim p}\|F_{2}(F_{1}(x))-x\|_{2}^{2}+\mathbb{E}_{y\sim q}\|F_{1}(F_{2}(y))-y\|_{2}^{2}.$
Vector space preservation (VSP) ensures that pairwise relationships between generated embeddings remain consistent under translation [40, 56]. Given a batch of $B$ embeddings $x_{1},...,x_{B}$ and $y_{1},...,y_{B}$, we sum their average pairwise distances after translation by both $F_{1}$ and $F_{2}$:
$\mathcal{L}_{\text{VSP}}(F_{1},F_{2})=\frac{1}{B}\sum_{i=1}^{B}\sum_{j=1}^{B}\bigg[$ $\|M_{1}(x_{i})\cdot M_{1}(x_{j})-F_{1}(M_{1}(x_{i}))\cdot F_{1}(M_{1}(x_{j}))\|_{2}^{2}$
$+\|M_{2}(y_{i})\cdot M_{2}(y_{j})-F_{2}(M_{2}(y_{i}))\cdot F_{2}(M_{2}(y_{j}))\|_{2}^{2}\bigg]$
Combining these losses yields: $\mathcal{L}_{\text{gen}}(\theta)=\lambda_{\text{rec}}\mathcal{L}_{\text{rec}}(R_{1},R_{2})+\lambda_{\text{CC}}\mathcal{L}_{\text{CC}}(F_{1},F_{2})+\lambda_{\text{VSP}}\mathcal{L}_{\text{VSP}}(F_{1},F_{2})$, where hyperparameters $\lambda_{\text{CC}}$, $\lambda_{\text{rec}}$, and $\lambda_{\text{VSP}}$ control relative importance.
4 Experimental setup
4.1 Preliminaries
Datasets. We use the Natural Questions (NQ) [23] dataset of user queries and Wikipedia-sourced answers for training (a $2$-million subset) and evaluation (a $65536$ subset). To evaluate information extraction, we use TweetTopic [2], a dataset of tweets multi-labeled by 19 topics; a random $8192$-record subset of Pseudo Re-identified MIMIC-III (MIMIC) [26], a pseudo re-identified version of the MIMIC dataset [18] of patient records multi-labeled by 2673 MedCAT [22] disease descriptions; and a random $50$-email subset of the Enron Email Corpus (Enron) [19], an unlabeled, public dataset of internal emails of a defunct energy company.
Models. Table˜1 lists the embedding models representing three size categories, four transformer backbones, and two output dimensionalities. Granite is multilingual, CLIP is multimodal.
Model Params (M) Backbone Year Dims Max Seq.
[41] gtr 110 T5 2021 768 512
[44] clip 151 CLIP 2021 512 77
[51] e5 109 BERT 2022 768 512
[29] gte 109 BERT 2023 768 512
[59] stella 109 BERT 2023 768 512
[13] granite 278 RoBERTa 2024 768 512
Table 1: Embedding models used in our experiments.
Training. Unless otherwise specified, each vec2vec is trained on two sets of embeddings generated from disjoint sets of $1$ million $64$-token sequences sampled from NQ (see Appendix˜C for experiments with fewer embeddings). Due to GAN instability [47], we select the best of multiple initializations and leave more robust training to future work.
4.2 Evaluating translation
Let $u_{i}=M_{1}(d_{i})$ and $v_{i}=M_{2}(d_{i})$ denote the source and target embeddings of the same input $d_{i}$. The goal of translation is to generate a vector that is as close to $v_{i}$ is possible. We say that $(u_{i},v_{j})$ are “aligned” by the translator $F$ if $v_{j}$ is the closest embedding to $F(u_{i})$: $j=\arg\min_{k}\cos\bigl(F(u_{i}),v_{k}\bigr).$ A perfect translator $F^{*}$ satisfies $i=\arg\min_{k}\cos\bigl(F^{*}(u_{i}),v_{k}\bigr)$ for all $i$.
Given (unknown) embeddings $\{M_{2}(d_{j})\}_{j=0}^{n}$ ordered by decreasing cosine similarity to $F(u_{i})$, let $r_{i}$ be the rank of the correct embedding $v_{i}=M_{2}(d_{i})$. To measure quality of $F$, we use three metrics. Mean Cosine Similarity measures how close translations are, on average, to their targets. Top-1 Accuracy is the fraction of translations whose target is closer than any other embedding. Mean Rank is the average rank of targets with respect to translations. The ideal translator $F^{*}$ achieves mean similarity of $1.0$, top-1 accuracy of $1.0$, and mean rank of $1.0$. Recall that a random alignment corresponds to a mean rank of $\frac{n}{2}$. Formally,
$\cos(u_{i},v_{i})=\frac{1}{n}\sum_{i=1}^{n}\bigl[1-\cos\bigl(F(u_{i}),v_{i}\bigr)\bigr]\quad\text{Top-1}(r)=\frac{1}{n}\sum_{i=1}^{n}\mathbf{1}\{r_{i}=1\}\quad\text{Rank(r)}=\frac{1}{n}\sum_{i=1}^{n}r_{i}$
vec2vec is the first unsupervised embedding translator, thus there is no direct baseline. As our Naïve baseline, we simply use $F(x)=x$ to measure geometric similarity between embedding spaces. The second (pseudo)baseline is Oracle-aided optimal assignment. It assumes that candidate targets are known and is thus strictly easier than vec2vec and the Naïve baseline. We solve optimal assignment, $\pi^{*}=\arg\min_{\pi}\sum_{i=1}^{n}\cos(u_{i},v_{\pi(i)})$, via either the Hungarian, Earth Mover’s Distance, Sinkhorn, or Gromov-Wasserstein algorithms, choosing the best-performing solver per experiment.
4.3 Evaluating information extraction
We measure whether translation preserves semantics via attribute inference: for each translated embedding $F(M_{1}(d_{i}))$, our goal is to infer attributes $c_{i}\subseteq\mathcal{C}$ of $d_{i}$.
The first method we use is zero-shot embedding attribute inference: calculate pairwise cosine similarities between $F(M_{1}(d_{i}))$ and the embeddings of all attributes in $\mathcal{C}$, identify top $k$ closest attributes, and measure whether they are correct via top-$k$ accuracy: $\frac{1}{n}\sum_{i=0}^{n}\mathbf{1}\left\{|c^{k}_{i}\cap c_{i}|\geq 1\right\}$.
The second method is embedding inversion that recovers text inputs from embeddings. Since [37] requires a pre-trained inversion model for each embedding space, we use [58] instead to generate an approximation $d^{\prime}_{i}$ of $d_{i}$ from $F(M_{1}(d_{i}))$ in a zero-shot manner. We measure the extracted information using LLM judge accuracy: the fraction of translated embeddings for which GPT-4o determines that $d^{\prime}$ reveals information in $d$. See Appendix˜D for our prompt.
In addition to the Naïve baseline, we also consider an Oracle attribute inference: zero-shot classification with the correct embedding $M_{2}(d)$ and class labels $M_{2}(\mathcal{C})$.
5 vec2vec learns to translate embeddings without any paired data
We first show that vec2vec learns a universal latent space, then demonstrate that this space preserves the geometry of all embeddings. Therefore, we can use it like a universal language of text encoders to translate their representations without any paired data.
Figure 4: Pairwise cosine similarities of input embeddings (left) and their vec2vec latents (middle) across different embedding pairs. The absolute difference between the heatmaps plots is on the right. All numbers are computed on the same batch of 1024 NQ texts.
vec2vec learns a universal latent space. vec2vec projects embeddings $M_{1,2,\ldots}$ into a shared latent space via compositions of input adapters ($A_{1,2,\ldots}$) and a shared translator $T$. Figure˜4 shows that even when the embeddings $u_{i}=M_{1}(d_{i})$ and $v_{i}=M_{2}(d_{i})$ are far apart (i.e., have low cosine similarity), their representations in vec2vec’s latent space are incredibly close: $T(A_{1}(u_{i}))\approx T(A_{2}(v_{i}))$. Figure˜1 visualizes this (via two-dimensional projections) for vec2vec trained on gte and gtr embeddings: the embeddings are far apart, but their latents are nearly overlapping.
vec2vec Naïve Baseline OA Baseline
$M_{1}$ $M_{2}$ $\cos(\cdot)$ $\uparrow$ Top-1 $\uparrow$ Rank $\downarrow$ $\cos(\cdot)$ $\uparrow$ Top-1 $\uparrow$ Rank $\downarrow$ Top-1 $\uparrow$ Rank $\downarrow$
gran. gtr 0.80 (0.0) 0.99 1.19 (0.1) -0.03 (0.0) 0.00 4168.73 (9.2) 0.00 4096.75 (9.2)
gte 0.87 (0.0) 0.95 1.18 (0.0) 0.01 (0.0) 0.00 4088.58 (9.2) 0.00 4069.91 (9.2)
stel. 0.79 (0.0) 0.98 1.05 (0.0) 0.01 (0.0) 0.00 4208.26 (9.2) 0.00 4096.75 (9.2)
e5 0.85 (0.0) 0.98 1.11 (0.0) 0.02 (0.0) 0.00 4111.60 (9.2) 0.00 4096.75 (9.2)
gtr gran. 0.81 (0.0) 0.99 1.02 (0.0) -0.03 (0.0) 0.00 4169.76 (9.2) 0.00 4096.66 (9.2)
gte 0.87 (0.0) 0.93 2.31 (0.1) 0.04 (0.0) 0.00 4080.92 (9.2) 0.00 4079.92 (9.2)
stel. 0.80 (0.0) 0.99 1.03 (0.0) 0.00 (0.0) 0.00 4198.78 (9.2) 0.00 4096.69 (9.2)
e5 0.83 (0.0) 0.84 2.88 (0.2) 0.03 (0.0) 0.00 4082.84 (9.2) 0.00 4066.42 (9.2)
gte gran. 0.75 (0.0) 0.95 1.22 (0.0) 0.01 (0.0) 0.00 4079.81 (9.3) 0.00 4069.23 (9.2)
gtr 0.75 (0.0) 0.91 2.64 (0.1) 0.04 (0.0) 0.00 4084.15 (9.2) 0.00 4078.45 (9.2)
stel. 0.89 (0.0) 1.00 1.00 (0.0) 0.56 (0.0) 1.00 1.00 (0.0) 1.00 1.00 (0.0)
e5 0.87 (0.0) 0.99 5.19 (0.5) 0.68 (0.0) 1.00 1.00 (0.0) 1.00 1.00 (0.0)
stel. gran. 0.80 (0.0) 0.98 1.08 (0.0) 0.01 (0.0) 0.00 4209.08 (9.3) 0.00 4096.88 (9.2)
gtr 0.82 (0.0) 1.00 1.10 (0.0) 0.00 (0.0) 0.00 4192.31 (9.2) 0.00 4096.63 (9.2)
gte 0.92 (0.0) 1.00 1.00 (0.0) 0.56 (0.0) 1.00 1.00 (0.0) 1.00 1.00 (0.0)
e5 0.86 (0.0) 1.00 1.00 (0.0) 0.38 (0.0) 0.99 1.03 (0.0) 1.00 1.00 (0.0)
e5 gran. 0.81 (0.0) 0.99 2.20 (0.2) 0.02 (0.0) 0.00 4120.60 (9.3) 0.00 4096.36 (9.2)
gtr 0.74 (0.0) 0.82 2.56 (0.0) 0.03 (0.0) 0.00 4080.76 (9.3) 0.00 4065.74 (9.2)
gte 0.90 (0.0) 1.00 1.01 (0.0) 0.68 (0.0) 1.00 1.00 (0.0) 1.00 1.00 (0.0)
stel. 0.78 (0.0) 1.00 1.00 (0.0) 0.38 (0.0) 1.00 1.00 (0.0) 1.00 1.00 (0.0)
Table 2: In-distribution translations: vec2vecs trained on NQ and evaluated on a 65536 text subset of NQ (chunked in batches of size 8192). The rank metric varies from 1 to 8192, thus 4096 corresponds to a random ordering. Since Optimal Assignment is a matching method, cosine distances are not applicable. Standard errors are shown in parentheses. Bold denotes best value.
vec2vec translations mirror target geometry. Table˜2 shows that vec2vec generates embeddings with near-optimal assignment across model pairs, achieving cosine similarity scores up to 0.92, top-1 accuracies up to 100%, and ranks as low as 1. In same-backbone pairings (e.g., (gte, e5)), vec2vec’s top-1 accuracy and rank are comparable to both the naïve baseline and (surprisingly) the oracle-aided optimal assignment. Although the embeddings generated by vec2vec are significantly closer to the ground truth than the naïve baseline, in same-backbone pairings the embeddings are close enough to be compatible. In cross-backbone pairings, vec2vec is far superior on all metrics, while baseline methods perform similarly to random guessing.
TweetTopic MIMIC
$M_{1}$ $M_{2}$ $\cos(\cdot)$ $\uparrow$ Top-1 $\uparrow$ Rank $\downarrow$ $\cos(\cdot)$ $\uparrow$ Top-1 $\uparrow$ Rank $\downarrow$
gran. gtr 0.74 (0.0) 0.99 1.09 (0.1) 0.74 (0.0) 0.60 23.38 (1.6)
gte 0.85 (0.0) 0.95 1.26 (0.1) 0.85 (0.0) 0.08 346.21 (7.8)
stel. 0.77 (0.0) 0.96 1.11 (0.0) 0.72 (0.0) 0.13 242.23 (6.1)
e5 0.83 (0.0) 0.87 3.10 (0.7) 0.84 (0.0) 0.12 361.06 (8.7)
gtr gran. 0.79 (0.0) 0.98 2.41 (0.6) 0.78 (0.0) 0.51 35.27 (1.9)
gte 0.85 (0.0) 0.96 1.29 (0.2) 0.84 (0.0) 0.12 279.56 (6.9)
stel. 0.77 (0.0) 0.96 1.10 (0.0) 0.72 (0.0) 0.27 127.92 (4.4)
e5 0.80 (0.0) 0.53 13.38 (1.2) 0.82 (0.0) 0.01 1413.80 (18.3)
gte gran. 0.73 (0.0) 0.94 1.33 (0.1) 0.73 (0.0) 0.09 342.15 (7.8)
gtr 0.71 (0.0) 0.95 1.29 (0.1) 0.69 (0.0) 0.12 256.63 (6.4)
stel. 0.86 (0.0) 1.00 1.00 (0.0) 0.85 (0.0) 1.00 1.00 (0.0)
e5 0.83 (0.0) 0.91 1.57 (0.2) 0.86 (0.0) 0.54 17.71 (0.9)
stel. gran. 0.79 (0.0) 0.99 1.09 (0.1) 0.77 (0.0) 0.14 221.95 (5.9)
gtr 0.77 (0.0) 1.00 1.00 (0.0) 0.75 (0.0) 0.56 17.70 (1.0)
gte 0.90 (0.0) 1.00 1.00 (0.0) 0.91 (0.0) 1.00 1.00 (0.0)
e5 0.85 (0.0) 0.98 1.05 (0.0) 0.85 (0.0) 0.51 26.33 (1.2)
e5 gran. 0.79 (0.0) 0.98 1.08 (0.0) 0.78 (0.0) 0.21 151.09 (4.6)
gtr 0.67 (0.0) 0.80 3.10 (0.6) 0.66 (0.0) 0.01 1029.64 (14.9)
gte 0.87 (0.0) 0.99 1.02 (0.0) 0.87 (0.0) 0.60 32.59 (2.6)
stel. 0.75 (0.0) 0.98 1.06 (0.0) 0.75 (0.0) 0.46 32.12 (1.4)
Table 3: Out-of-distribution translations: vec2vecs trained on NQ and evaluated on the entire TweetTopic test set (800 tweets) and an 8192-record subset of MIMIC. The rank metric varies from 1 to 800 (for TweetTopic) and 8192 (for MIMIC), thus 400 and, respectively, 4096 correspond to a random ordering. Standard errors are shown in parentheses.
Table˜3 shows that this performance extends to out-of-distribution data. Our vec2vec translators were trained on NQ (drawn from Wikipedia), yet exhibit high cosine similarity, high accuracy, and low rank when evaluated on tweets (which are far more colloquial and use emojis) and medical records (which contain domain-specific jargon unlikely to appear in NQ). In Appendix˜B, we show that baseline methods fail on cross-backbone embedding pairs.
vec2vec OA Baseline
$M_{1}$ $M_{2}$ $\cos(\cdot)$ $\uparrow$ Top-1 $\uparrow$ Rank $\downarrow$ Top-1 $\uparrow$ Rank $\downarrow$
gran. clip 0.78 (0.0) 0.35 226.62 (3.2) 0.00 4096.32 (9.2)
gtr 0.73 (0.0) 0.13 711.23 (5.9) 0.00 4096.71 (9.2)
gte 0.62 (0.0) 0.00 3233.41 (9.8) 0.00 4096.32 (9.2)
stel. 0.77 (0.0) 0.31 286.69 (3.6) 0.00 4096.60 (9.2)
e5 0.64 (0.0) 0.01 2568.21 (9.4) 0.00 4096.78 (9.2)
clip gran. 0.74 (0.0) 0.72 4.46 (0.1) 0.00 4096.36 (9.2)
gtr 0.67 (0.0) 0.27 155.11 (2.1) 0.00 4096.80 (9.2)
gte 0.75 (0.0) 0.00 2678.90 (8.9) 0.00 4096.12 (9.2)
stel. 0.72 (0.0) 0.61 22.50 (0.5) 0.00 4096.47 (9.2)
e5 0.73 (0.0) 0.01 1692.28 (8.2) 0.00 4096.44 (9.2)
Table 4: Translations between unimodal and multimodal (CLIP) embeddings: vec2vecs trained on NQ and evaluated on a 65536 text subset of NQ (chunked in batches of size 8192). Rank varies from 1 to 8192, thus 4096 corresponds to a random ordering. Since the embedding dimensionalities are different, the naive baseline does not apply. Bold denotes best value.
Finally, Table˜4 shows that vec2vec can even translate to and from the space of CLIP, a multimodal embedding model which was trained in part on image data. While the translations are not as strong as in Table˜2, vec2vec consistently outperforms the optimal assignment baseline. These results show that promise of our method at adapting to new modalities: in particular, the embedding space of CLIP has been successfully connected to other modalities such as heatmaps, audio, and depth charts [11].
6 Using vec2vec translations to extract information
In this section, we show that vec2vec translations not only preserve the geometric structure of embeddings but also retain sufficient semantics to enable attribute inference.
TweetTopic ($k=1$) MIMIC ($k=10$)
$M_{1}$ $M_{2}$ vec2vec Naïve $M_{1}$ $M_{2}$ vec2vec Naïve $M_{1}$ $M_{2}$
gran. gtr 0.25 0.10 0.30 0.24 0.19 0.11 0.76 0.88
gte 0.32 0.09 0.30 0.34 0.36 0.13 0.76 1.00
stel. 0.24 0.10 0.30 0.28 0.27 0.04 0.76 0.96
e5 0.31 0.18 0.30 0.31 0.19 0.20 0.76 0.97
gtr gran. 0.34 0.08 0.24 0.30 0.16 0.12 0.88 0.76
gte 0.33 0.13 0.24 0.34 0.28 0.05 0.88 1.00
stel. 0.30 0.10 0.24 0.28 0.25 0.07 0.88 0.96
e5 0.30 0.04 0.24 0.31 0.09 0.09 0.88 0.97
gte gran. 0.37 0.04 0.34 0.30 0.18 0.11 1.00 0.76
gtr 0.24 0.13 0.34 0.24 0.10 0.03 1.00 0.88
stel. 0.31 0.20 0.34 0.28 0.68 0.83 1.00 0.96
e5 0.37 0.30 0.34 0.31 0.37 0.63 1.00 0.97
stel. gran. 0.35 0.07 0.28 0.30 0.23 0.09 0.96 0.76
gtr 0.26 0.13 0.28 0.24 0.22 0.09 0.96 0.88
gte 0.38 0.36 0.28 0.34 0.90 0.98 0.96 1.00
e5 0.35 0.34 0.28 0.31 0.38 0.46 0.96 0.97
e5 gran. 0.33 0.15 0.31 0.30 0.14 0.07 0.97 0.76
gtr 0.26 0.22 0.31 0.24 0.11 0.04 0.97 0.88
gte 0.34 0.28 0.31 0.34 0.47 0.66 0.97 1.00
stel. 0.26 0.16 0.31 0.28 0.36 0.40 0.97 0.96
Table 5: Information leakage via top-$k$ zero-shot attribute inference: vec2vecs trained on NQ and evaluated on the TweetTopic test set (800 tweets) and an 8192-record subset of MIMIC. $M_{1}$ and $M_{2}$ represent ideal zero-shot inference: attributes and embeddings are encoded using the same model.
Figure 5: Leakage of information via inversion. Trained on NQ and evaluated on a 50-email subset of the Enron Email Corpus. Cells denote judge accuracy.
Zero-shot attribute inference. Table˜5 shows that attribute inference on vec2vec translations consistently outperforms the naïve baseline and often does better than the ideal zero-shot baseline which performs inference on ground-truth document and attribute embeddings in the same space (this baseline is imaginary since these embeddings are not available in our setting).
vec2vec translations even work for embeddings of medical records, which are much further from the training distribution than tweets. The attributes in this case are MedCAT disease descriptions, very few of which occur in the training data. Attribute inference on translated embeddings is comparable to the naïve baseline in same-backbone pairings and outperforms it (often greatly) in cross-backbone pairings. The fact that vec2vec preserves the semantics of concepts like "alveolar periostitis" (which never appears in its training data) is evidence that its latent space is indeed a universal representation.
Zero-shot inversion. Inversion, i.e., reconstruction of text inputs, is more ambitious than attribute inference. vec2vec translations retain enough semantic information that off-the-shelf, zero-shot inversion methods like [58], developed for embeddings computed by standard encoders, extract information for as many as 80% of documents given only their translated embeddings, for some model pairs (Figure˜5). These inversions are imperfect and we leave development of specialized inverters for translated embeddings to future work. Nevertheless, as exemplified in Figure˜6, they still extract information such as individual and company names, dates, promotions, financial information, outages, and even lunch orders. In Appendix˜D, we show the prompt we use to measure extraction.
[…]
Figure 6: Examples of inversions that infer entities and content.
7 Related work
Representation alignment. Similarities between representations of different neural networks are investigated in [24, 28, 52, 4, 17]. CCA [36], SVCCA, [45], CKA [21], and GUI-based systems such as embComp [15] compare embeddings from different subspaces. Methods like [33, 39, 34, 42] harness representation similarity for zero-shot stitching, substitution, and multi-modal adaptation. All rely on some amount of paired data, which is difficult to reduce [5]. Our method does not just measure similarity, we learn how to translate representations across spaces without any paired data.
Unsupervised transport. The problem of unsupervised transport has been studied for images [16, 32, 60], words [54, 9, 14, 8], and natural language sequences [46, 25, 1, 3, 55]. Our method builds on these works, which often employ a combination of cycle-consistency and adversarial loss. [48] proposes a solver for matching small sets of embeddings between different vision-language models. Our method goes well beyond matching by taking unknown embeddings and generating matching embeddings in the space of another model.
Embedding inversion. An emerging line of research investigates decoding text from language model embeddings [49, 27, 37] and outputs [38, 6, 57]. vec2vec helps apply these to unknown embeddings, without an encoder or paired data, by translating them to the space of a known model.
Bridging modality gaps. Previous work has noted an inherent “gap” between image- and text-based models [30] and proposed various ways to unify the modalities [50]. Some approaches feed image embeddings directly into language models [20, 53, 10, 31], while others generate captions from image embeddings [35] or even from text embeddings themselves [37]. [11] introduces a shared embedding space that integrates inputs from multiple modalities, including text, audio, and vision. In contrast, our post-hoc approach directly translates between representations and complements these systems by enabling inputs from a wide variety of embedding models.
8 Discussion and Future Work
The Platonic Representation Hypothesis conjectures that the representation spaces of modern neural networks are converging. We assert the Strong Platonic Representation Hypothesis: the latent universal representation can be learned and harnessed to translate between representation spaces without any encoders or paired data.
In Section˜5, we demonstrated that our vec2vec method successfully translates embeddings generated from unseen documents by unseen encoders, and the translator is robust to (sometimes very) out-of-distribution inputs. This suggests that vec2vec learns domain-agnostic translations based on the universal geometric relationships which encode the same semantics in multiple embedding spaces.
In Section˜6, we showed that vec2vec translations preserve sufficient input semantics to enable attribute inference. We extracted sensitive disease information from patient records and partial content from corporate emails, with access only to document embeddings and no access to the encoder that produced them. Better translation methods will enable higher-fidelity extraction, confirming once again that embeddings reveal (almost) as much as their inputs.
Our findings provide compelling evidence for the Strong Platonic Representation Hypothesis for text-based models. Our preliminary results on CLIP suggest that the universal geometry can be harnessed in other modalities, too. The results in this paper are but a lower bound on inter-representation translation. Better and more stable learning algorithms, architectures, and other methodological improvements will support scaling to more data, more model families, and more modalities.
Acknowledgments and Disclosure of Funding
This research is supported in part by the Google Cyber NYC Institutional Research Program. JM is supported by the National Science Foundation.
Appendix A Compute
Our training and evaluation were conducted using diverse compute environments, including both local and cloud GPU clusters. Experiments were done on NVIDIA 2080Ti, L4, A40, and A100 GPUs, listed in order of increasing computational capacity.
For our main experiments, we trained 36 distinct vec2vec models, with training durations ranging from 1 to 7 days per model, depending on the specific GPU and model pair (which affected convergence rates). Additional 9 vec2vec models were trained for our ablations, bringing the total to 45 vec2vec models. Taking a conservative estimate of the average training time, this amounted to approximately 180 GPU days (45 models $\times$ 4 days / model).
Evaluation procedures varied by model type:
  • For each of our 45 vec2vec models, our full evaluation on NQ, TweetTopic, and MIMIC datasets required approximately 1 hour per model across all GPU types.
  • The 36 non-ablation GANs underwent additional analysis on the Enron email corpus, requiring about 1.5 hours for inversion and downstream LLM evaluation.
  • Naive baselines (30 models without CLIP) required approximately 30 minutes each for evaluation, roughly half the time needed for our full models.
  • Optimal assignment baselines (36 models) were run exclusively on CPU. We implemented 4 different assignment algorithms, each requiring approximately 4 hours of computation time. The Gromov-Wasserstein algorithm was particularly computationally intensive and was included for CLIP due to dimensional differences.
In total, our experiments consumed approximately 180 GPU days for training and an additional 78 GPU hours for evaluation and analysis. There were approximately 144 CPU hours required for optimal assignment.
Appendix B Full out-of-distribution translation results
We provide baseline numbers for the experiments shown in Table˜3, by dataset.
vec2vec Naïve Optimal Transport
$E_{1}$ $E_{2}$ $\cos(\cdot)$ Top-1 Rank $\cos(\cdot)$ Top-1 Rank Top-1 Rank
gte gtr 0.71 (0.0) 0.95 1.29 (0.1) 0.04 (0.0) 0.00 386.58 (8.3) 0.00 383.41 (8.1)
stel. 0.86 (0.0) 1.00 1.00 (0.0) 0.58 (0.0) 1.00 1.00 (0.0) 1.00 1.00 (0.0)
e5 0.83 (0.0) 0.91 1.57 (0.2) 0.68 (0.0) 1.00 1.00 (0.0) 1.00 1.00 (0.0)
gran. 0.73 (0.0) 0.94 1.33 (0.1) 0.00 (0.0) 0.00 408.81 (8.3) 0.01 398.16 (8.2)
gtr gte 0.85 (0.0) 0.96 1.29 (0.2) 0.04 (0.0) 0.00 392.01 (8.2) 0.00 384.00 (8.1)
stel. 0.77 (0.0) 0.96 1.10 (0.0) 0.00 (0.0) 0.00 394.69 (8.3) 0.00 390.67 (8.2)
e5 0.80 (0.0) 0.53 13.38 (1.2) 0.03 (0.0) 0.00 400.85 (8.2) 0.00 400.48 (8.2)
gran. 0.79 (0.0) 0.98 2.41 (0.6) -0.04 (0.0) 0.00 411.53 (8.3) 0.00 401.01 (8.2)
stel. gte 0.90 (0.0) 1.00 1.00 (0.0) 0.58 (0.0) 1.00 1.00 (0.0) 1.00 1.00 (0.0)
gtr 0.77 (0.0) 1.00 1.00 (0.0) 0.00 (0.0) 0.00 393.07 (8.1) 0.00 390.26 (8.2)
e5 0.85 (0.0) 0.98 1.05 (0.0) 0.37 (0.0) 0.89 1.55 (0.1) 1.00 1.00 (0.0)
gran. 0.79 (0.0) 0.99 1.09 (0.1) 0.00 (0.0) 0.00 418.16 (8.4) 0.00 400.95 (8.2)
e5 gte 0.87 (0.0) 0.99 1.02 (0.0) 0.68 (0.0) 1.00 1.00 (0.0) 1.00 1.00 (0.0)
gtr 0.67 (0.0) 0.80 3.10 (0.6) 0.03 (0.0) 0.00 401.16 (8.4) 0.00 400.49 (8.2)
stel. 0.75 (0.0) 0.98 1.06 (0.0) 0.37 (0.0) 1.00 1.00 (0.0) 1.00 1.00 (0.0)
gran. 0.79 (0.0) 0.98 1.08 (0.0) 0.02 (0.0) 0.00 405.75 (8.3) 0.00 401.00 (8.2)
gran. gte 0.85 (0.0) 0.95 1.26 (0.1) 0.00 (0.0) 0.00 406.73 (8.2) 0.01 398.14 (8.2)
gtr 0.74 (0.0) 0.99 1.09 (0.1) -0.04 (0.0) 0.00 415.61 (8.2) 0.00 400.99 (8.2)
stel. 0.77 (0.0) 0.96 1.11 (0.0) 0.00 (0.0) 0.00 417.27 (8.2) 0.00 400.97 (8.2)
e5 0.83 (0.0) 0.87 3.10 (0.7) 0.02 (0.0) 0.00 405.53 (8.1) 0.00 401.00 (8.2)
Table 6: Out-of-distribution translations on TweetTopic (with baselines): vec2vec models trained on NQ and evaluated on the entire TweetTopic test set (800 tweets). The rank metric varies from 1 to 800, thus 400 corresponds to a random ordering. Since Optimal Assignment is a matching method, cosine distances are not applicable. Standard errors are shown in parentheses.
vec2vec Naïve Optimal Transport
$E_{1}$ $E_{2}$ $\cos(\cdot)$ Top-1 Rank $\cos(\cdot)$ Top-1 Rank Top-1 Rank
gte gtr 0.69 (0.0) 0.12 256.63 (6.4) 0.08 (0.0) 0.00 4229.90 (26.2) 0.00 4096.63 (26.1)
stel. 0.85 (0.0) 1.00 1.00 (0.0) 0.56 (0.0) 1.00 1.00 (0.0) 1.00 1.00 (0.0)
e5 0.86 (0.0) 0.54 17.71 (0.9) 0.69 (0.0) 0.98 1.04 (0.0) 1.00 1.00 (0.0)
gran. 0.73 (0.0) 0.09 342.15 (7.8) 0.01 (0.0) 0.00 3946.19 (25.8) 0.00 3802.92 (25.9)
gtr gte 0.84 (0.0) 0.12 279.56 (6.9) 0.08 (0.0) 0.00 4180.47 (26.2) 0.00 4096.99 (26.1)
stel. 0.72 (0.0) 0.27 127.92 (4.4) 0.00 (0.0) 0.00 4296.04 (26.1) 0.00 4096.53 (26.1)
e5 0.82 (0.0) 0.01 1413.80 (18.3) 0.09 (0.0) 0.00 4064.47 (26.2) 0.00 4010.13 (26.1)
gran. 0.78 (0.0) 0.51 35.27 (1.9) -0.02 (0.0) 0.00 4023.67 (26.1) 0.00 3964.83 (26.1)
stel. gte 0.91 (0.0) 1.00 1.00 (0.0) 0.56 (0.0) 1.00 1.00 (0.0) 1.00 1.00 (0.0)
gtr 0.75 (0.0) 0.56 17.70 (1.0) 0.00 (0.0) 0.00 4339.83 (26.2) 0.00 4097.00 (26.1)
e5 0.85 (0.0) 0.51 26.33 (1.2) 0.35 (0.0) 0.59 12.68 (0.6) 1.00 1.00 (0.0)
gran. 0.77 (0.0) 0.14 221.95 (5.9) -0.01 (0.0) 0.00 3951.42 (25.9) 0.00 3776.52 (26.0)
e5 gte 0.87 (0.0) 0.60 32.59 (2.6) 0.69 (0.0) 0.98 1.09 (0.0) 1.00 1.00 (0.0)
gtr 0.66 (0.0) 0.01 1029.64 (14.9) 0.09 (0.0) 0.00 4032.85 (26.2) 0.00 4010.06 (26.1)
stel. 0.75 (0.0) 0.46 32.12 (1.4) 0.35 (0.0) 0.86 2.49 (0.1) 1.00 1.01 (0.0)
gran. 0.78 (0.0) 0.21 151.09 (4.6) 0.02 (0.0) 0.00 4008.10 (25.9) 0.00 3932.58 (26.2)
gran. gte 0.85 (0.0) 0.08 346.21 (7.8) 0.01 (0.0) 0.00 3978.35 (26.1) 0.00 3808.18 (25.9)
gtr 0.74 (0.0) 0.60 23.38 (1.6) -0.02 (0.0) 0.00 4010.00 (25.8) 0.00 3962.83 (26.1)
stel. 0.72 (0.0) 0.13 242.23 (6.1) -0.01 (0.0) 0.00 3900.74 (26.2) 0.00 3780.44 (26.0)
e5 0.84 (0.0) 0.12 361.06 (8.7) 0.02 (0.0) 0.00 4024.92 (26.1) 0.00 3937.63 (26.2)
Table 7: Out-of-distribution translations on MIMIC (with baselines): vec2vec models trained on NQ and evaluated on an 8192-record subset of MIMIC. The rank metric varies from 1 to 8192, thus 4096 corresponds to a random ordering. Since Optimal Assignment is a matching method, cosine distances are not applicable. Standard errors are shown in parentheses.
Appendix C Ablations
C.1 Ablating components of our method
We ablate our method subtractively, measuring the key metrics after removing individual components of our algorithm (described in Section˜3). Table˜8 shows that each component appears to be critical to building good translations. While in each setting, vec2vec’s $\cos(\cdot)$ is higher than the naïve baselines, the Top-1 accuracies and ranks imply that ablated vec2vec translations are at best only slightly better than the baseline and do not preserve the geometry of the vector space.
Method $\cos(\cdot)$ Top-1 Rank
vec2vec 0.75 (0.0) 0.91 2.64 (0.1)
Naïve 0.04 (0.0) 0.00 4084.15 (9.2)
Optimal Transport - 0.00 4078.45 (9.2)
– VSP loss 0.58 (0.0) 0.00 4196.64 (9.2)
– CC loss 0.50 (0.0) 0.00 3941.36 (9.3)
– latent GAN 0.49 (0.0) 0.00 3897.09 (9.5)
– VSP and CC loss 0.47 (0.0) 0.00 3365.24 (9.3)
– hyperparam. tuning 0.50 (0.0) 0.00 4011.73 (9.3)
Table 8: gte $\to$ gtr translators trained without individual components of our method on NQ and evaluated on a 65536-text subset of NQ (chunked in batches of 8192). The rank metric varies from 1 to 8192, thus 4096 corresponds to a random ordering. Since Optimal Assignment is a matching method, cosine distances are not applicable. Standard errors are shown in parentheses.
C.2 Amount of data needed to learn translation
NN $\cos(\cdot)$ Top-1 Rank
1000000 0.75 (0.0) 0.92 2.73 (0.2)
10000 0.57 (0.0) 0.01 1462.21 (20.)
50000 0.74 (0.0) 0.81 3.91 (0.6)
100000 0.74 (0.0) 0.85 4.52 (0.4)
500000 0.75 (0.0) 0.92 2.73 (0.2)
Table 9: gte $\to$ gtr translators trained with different amounts of gte data: vec2vec models trained on NQ and evaluated an 8192-record subset of NQ. The rank metric varies from 1 to 8192, thus 4096 corresponds to a random ordering. Since Optimal Assignment is a matching method, cosine distances are not applicable. Standard errors are shown in parentheses.
In our main experiments, we use 1M-point subsets of NQ to train our vec2vec models. Now, we train the gtr-gte vec2vec with 1M gtr embeddings but fewer gte embeddings. Table˜9 shows that translators still learn something (i.e. are better than random), and only 50K embeddings are sufficient to learn a translation that is almost as good. Even with 10K embeddings, our method still learns a non-trivial translation. Translation performance generally seems to improve with the amount of training data, however.
Appendix D Prompt for measuring information extraction
Figure˜5 utilizes an LLM judge to determine whether the approximate inversion (using [58]) of a translated email embedding leaks information about the original email. Figure˜7 shows our prompt to the GPT-4o judge:
Original email: {ground_truth} Reconstructed email: {generation}. Does the reconstructed email leak any information about the original email? Answer with only ‘yes’ or ‘no’.
Figure 7: The prompt given to the LLM judge.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: Our main claim is that we are (1) able to translate embeddings with high cosine similarity across model pairs with different architectures, parameter counts, and training datasets and (2) these translations can be used for attribute inference. These are the two experimental sections exactly.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: The key limitation of our method is the training stability of GANs. Downstream of this limitation is long computation times and limited (while still effective) translation quality. This is discussed at length in the discussion.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [N/A]
  14. Justification: We do not show any theoretical results.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: Along with our releasing our code (used for each of the experiments we present), we describe in depth the loss functions and architecture choices we made. We also ablate these choices in the appendix.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: We link to well-documented code that will help reproduce any of our experiments.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: All details necessary to reconstruct our experiments are contained in our code. On top of that we describe our method at length in the body of our paper.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: Wherever applicable we show error bars: the rank and cosine metrics, namely.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: We include compute information in the appendix.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: All of our experiments are on publicly-available datasets that are either fully anonymous or pseudonymized.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: On the positive side, we provide evidence for a popular hypothesis in the ML community. On the negative side, we highlight, discuss, and provide remediation ideas for an emerging threat. As with many security-flavored papers, we hope to increase visibility of emerging threats.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: Although we have none planned, future releases of model weights would be of translators between known, open-source embedding spaces. These would not be very useful in exfiltrating data from unknown embeddings (a concern we highlight).
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: All data we use is publicly available, cited, and used as dictated by the licenses and terms.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [Yes]
  64. Justification: Our public code release will include documentation on how exactly to reconstruct our experiments.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: We did not use crowdsourcing or human subjects.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: We did not use crowdsourcing or human subjects.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [N/A]
  79. Justification: Our paper’s methodology does not involve LLMs as any important, original, or non-standard components.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.